OpenStudy (anonymous):

ok how do i get x if theres only 90 degrees?

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

there asking me to find x and the is only 90 degrees on gthediagram so how do i find x cus i have no freaking clue

OpenStudy (anonymous):

What is x? The distance from the corner to the point on the circle?

OpenStudy (anonymous):

270 90 = 1/2 (x- (360-x))

OpenStudy (anonymous):

same thing for 70 too like find x

OpenStudy (anonymous):

how ami supposed to find x if theres only just a number there like come on

OpenStudy (anonymous):

exterior angle formed by two tangents or secants equals 1/2 the difference of the major arc and minor arc

OpenStudy (radar):

90=1/2(x-(360-x)) 90=1/2x-1/2(360-x) 90=1/2x-180+1/2x 90=x-180 x=270 If mrjoselson is correct

OpenStudy (anonymous):

yes he is now i gotta find x for 70 still same set up

OpenStudy (radar):

Did you get x=250 degrees?

OpenStudy (anonymous):

at first i thot that x would still be 70

OpenStudy (radar):

I think x is the arc formed by the legs intersecting the circle, then angle is now 70 rather than 90 if it is the same "set up"

OpenStudy (anonymous):

no i mean do i need to dothe same for 90 like have 70 = 1/2(360-x)

OpenStudy (radar):

I am assuming so. Same formula using 70 rather than 90. that is how I came up with 250.

OpenStudy (anonymous):

ohhhh ok